OpenStudy (anonymous):

Solve the problem. Instruments on a satellite measure the amount of power generated by the satellite's power supply. The time t and the power P can be modeled by the function P = 50e^{-t/300}, where t is in days and P is in watts. How much power will be available after 378 days? Round to the nearest hundredth. a. 141.8 watts b. 14.18 watts c. 1.41 watts d. 0.14 watts

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

25% chance :)

OpenStudy (anonymous):

no im not

OpenStudy (anonymous):

Why not just plug in t = 378 in the expression?

OpenStudy (anonymous):

girl : Am I pretty? Boy : NO. Girl : Do you want to be with me forever? Boy : NO. Girl : Would you cry if I walked away?Boy : NO. She heard enough, and was hurt. She walked away, tears ran down her face. The boy grabbed her arm. Boy : Your not pretty, your beautiful. I don't want to be with you forever, I NEED to be with you forever. And I wouldn't cry if you walked away, I would DIE. (Boy whispers) : Please? Stay with me. (Girl whispers) : I will. Tonight at midnight your true love will realize they love you. Something good will happen to you between 1-4 pm. Tomorrow it could be anywhere. Get ready for the biggest shock of your life! If you don't post this to 5 other pages. You will have relationship problems for the next 10 years

OpenStudy (anonymous):

Jajaja you are funy...

OpenStudy (anonymous):

Plugging in t = 378 gives 14.18 watts, option b. If I have understodd the problem, this is correct.

OpenStudy (anonymous):

Elekt how can be sure that this the right answer?

OpenStudy (anonymous):

Well, if I have understood the problem, it asks how much power it is after 378 days. Since you have a function that describes the amount of power at any instant t, it should just be to plug in t = 378 days when t is in days :)

OpenStudy (anonymous):

ok you are right... thanks... i have mor question dear can you keep helping me ?

OpenStudy (anonymous):

Yes, sure :)

OpenStudy (anonymous):

List all the elements of B that belong to the given set. \[B =\left\{ 5,\sqrt{7,-3.0,} \right\} 0>5 0.62\] Rational numbers a. \{5,0\} b. \{\sqrt{7},\frac{0}{5},0.62\} c. \{ \sqrt{7} \} d. \{5,-3,0,\frac{0}{5},0.62\}

OpenStudy (anonymous):

Hmmz, not even sure what that means because of the \[0 > 50.62\] in there :P

OpenStudy (anonymous):

And are you just supposed to list the rational elements of that set?

OpenStudy (anonymous):